OpenStudy (anonymous):

help solving using substitution?? 2x+4y=3 and 4x+4y=8

OpenStudy (anonymous):

Next time try asking this in the math section. Substitution method is you have to substitute either one of the equation into one of the unknown. Let say; 1. 2x + 4y =3 2. 4x + 4y = 8 3. Let use (1) as a y-substitute equation; y = (3-2x)/4 4. Insert (3) into (2); 4x + (4(3-2x))/4 = 8, Find value of x. 5. Substitute value of x from (4) into (3) to find y.